OpenStudy (anonymous):

Which equation can pair with x + 2y = 5 to create an inconsistent system?

OpenStudy (anonymous):

here are the answer choices 2x+4y=3 5x+2y=3 6x+12y=30 3x+4y=8 i know its not the third one

OpenStudy (anonymous):

I think it's the first one :